Grade 12th passMechanics

a body is projected vertically upward from top of a tower with a velocity of 40 m per second .Tower being 180 m and after t second another body B allowed to fall from same point. Both bodies reach the ground simultaneously .calculate t and velocities of A and B on reaching the grounda

Sachin mishra
7 Years agoGrade 12th pass

1 Answer

Devang
7 Years ago
 let assume that object A throw vertically upward with velocity of   V0 = 40 m/s on the roof of the tower.
hight of tower H = 180m
Maximum hight attained by object = V02 / 2g             
                                                      =  40 * 40 / 2 * 10                                            (asume g= 10m/s2 )
                                             hmax = 80m
time taken by the object to reach maximum hight is = V0 / g
                                                                                  = 40 / 10 
                                                                             T1  = 4 sec.
 
 
time taken by object to reach the ground from maximum hight = (2 * h / g) ½       (where h = H + hmax
                                                                                                                                   = 180+80 = 260)
                                                                                                = (2 * 260 / 10)½                                         .                                                                                       T2  =  7.21 sec
total time to reach the ground is TA = T1 + T2
                                                      = 7.21 + 4
                                                 TA = 11.21 sec                           …...... (1)
velocity when object reach the ground = ( 2 *g *h)½   
                                                             = (2 *10 *260)½  
                                                       VA  = 72.11 m/s                   
let assume that onather object B is fall from top of the tower.
time taken by object B to reach the ground =   (2 * H / g) ½ 
                                                                    =   (2 * 180 / 10)½ 
                                                            TB   = 6 sec                   ….......(2)
velocity of object B when it touches the ground VB = ( 2 *g *H)½    
                                                                                = (2 *10 *180)½ 
                                                                            VB= 60 m
thrown upward object A and , after time t sec object B fall down 
and both object reaches the ground at same time
then     t = TA  –  TB                                              from (1) & (2) 
              = 11.21 – 6
           t  = 5.21 sec